# Ultra-Planckian Quark and Gluon Scattering in Agravity

### v1

@@ Abstract
We investigate tree-level scattering processes involving quarks and gluons mediated by graviton exchange in the framework of Agravity, a dimensionless and renormalizable theory of quadratic quantum gravity. Focusing on the ultra-Planckian regime, we compute the squared amplitudes and analyze the differential cross sections for the processes $gg\rightarrow gg$, $gg\rightarrow q\bar{q}$, $gq\rightarrow gq$, and $qq\rightarrow qq$. We demonstrate that all amplitudes scale as $1/s$ at high energies, in agreement with expectations for a UV-complete theory of gravity. In addition, we explore the issue of unitarity in the presence of higher-derivative ghost modes by analyzing the positivity properties of the squared amplitudes. While IR divergences appear in the forward scattering of massless particles, we show that these are regularized by finite quark masses. Our findings support the viability of Agravity as a perturbatively unitary and UV-complete extension of general relativity, capable of consistently describing gravitational interactions among elementary matter fields at trans-Planckian energies.
@@ I Introduction
In this context, quadratic quantum gravity (QQG) has been extensively studied as a promising candidate for a renormalizable theory of quantum gravity Stelle:1976gc ; Tomboulis:1977jk ; Odintsov:1991nd ; Salvio:2014soa ; Salvio:2017qkx ; Einhorn:2014gfa ; Buchbinder:2017lnd ; Donoghue:2021cza . Although such theories are renormalizable, they typically suffer from issues related to the presence of ghost-like or tachyonic modes. Recent investigations Salvio:2014soa ; Alvarez-Gaume:2015rwa ; Holdom:2015kbf ; Anselmi:2018ibi ; Salvio:2018crh ; Donoghue:2021cza ; Buccio:2024hys suggest that while ghost states might be controllable under certain prescriptions, the presence of tachyonic instabilities is generally regarded as a fundamental obstacle.
@@
Although Agravity is free from tachyonic instabilities, it is not, in principle, exempt from the presence of ghost-like states. This naturally raises the question: is it possible for Agravity to preserve unitarity despite the existence of ghosts? One avenue to explore this issue is through the explicit computation of scattering amplitudes mediated by graviton exchange within the framework of Agravity, with particular attention to the positivity properties of the squared amplitude. Furthermore, it is instructive to analyze the high-energy (ultra-Planckian) behavior of these amplitudes in an effort to determine whether the resulting differential cross sections remain physically meaningful in the deep UV regime.
Progress in this direction has been made by Holdom in Refs. Holdom:2021oii ; Holdom:2021hlo . In Ref. Holdom:2021oii , the author demonstrates that QQG leads to a well-behaved differential cross section for photon-photon scattering across all energy scales. The study also highlights how unitarity may be preserved even in the absence of positivity, with significant implications for the structure of scattering amplitudes and cross sections. In a subsequent work, Ref. Holdom:2021hlo , it is shown that high-energy scattering cross sections in quadratic gravity exhibit a $1/E^{2}$ scaling due to remarkable cancellations among the contributions of different graviton modes. This behavior provides further support for the idea that QQG may constitute a viable UV-complete theory of quantum gravity, capable of maintaining unitarity without relying on the positivity of individual amplitudes.
@@
The main objective of the present work is to compute graviton-mediated scattering amplitudes for quarks and gluons within the framework of QQG, specifically in the ultra-Planckian regime. To this end, we couple massless QCD fields to Agravity—a dimensionless and renormalizable model of QQG—and analyze the positivity of the squared scattering amplitudes, as well as the behavior of the corresponding differential cross sections in the deep UV.
@@ III Agravity-Mediated Scattering Processes
$\mathcal{M}_{t}$ $=$ $\mathcal{M}_{s}\Big{|}_{s\leftrightarrow t,\leavevmode\nobreak\ p_{3}\leftrightarrow-p_{2},\leavevmode\nobreak\ b\leftrightarrow c},$ (10)
$\mathcal{M}_{u}$ $=$ $\mathcal{M}_{s}\Big{|}_{u\leftrightarrow t,\leavevmode\nobreak\ p_{3}\leftrightarrow p_{4},\leavevmode\nobreak\ c\leftrightarrow d}.$ (11)
@@
$|\mathcal{M}|^{2}$ $=$ $\frac{f_{0}^{4}\left(N_{c}^{2}-1\right)\left[\left(2805N_{c}^{2}-448\right)tu-120t^{2}-120u^{2}\right]}{3tu}$ (12)
$-\frac{f_{0}^{2}f_{2}^{2}\left(N_{c}^{2}-1\right)}{6t^{2}u^{2}}\Big{[}2\left(117N_{c}^{2}-43\right)t^{4}+3\left(102N_{c}^{2}-155\right)t^{3}u$
$+2\left(75N_{c}^{2}-137\right)t^{2}u^{2}+3\left(102N_{c}^{2}-155\right)tu^{3}+2\left(117N_{c}^{2}-43\right)u^{4}\Big{]}$
$-\frac{f_{2}^{4}\left(N_{c}^{2}-1\right)}{3s^{2}t^{3}u^{3}}\Big{[}12\left(N_{c}^{2}-1\right)t^{8}+11\left(3N_{c}^{2}-4\right)t^{7}u+\left(33N_{c}^{2}-79\right)t^{6}u^{2}$
$+\left(18N_{c}^{2}-151\right)t^{5}u^{3}-196t^{4}u^{4}+\left(18N_{c}^{2}-151\right)t^{3}u^{5}+\left(33N_{c}^{2}-79\right)t^{2}u^{6}$
$+11\left(3N_{c}^{2}-4\right)tu^{7}+12\left(N_{c}^{2}-1\right)u^{8}\Big{]},$
where $N_{c}$ denotes the number of colors, with $N_{c}=3$ for QCD. The expression for $|\mathcal{M}|^{2}$ shown above corresponds to the squared matrix element fully summed over the color and polarization degrees of freedom of both the initial- and final-state gluons.
It is important to note that, similarly to the ultra-Planckian scattering of photons Holdom:2021oii , the differential cross section for gluon scattering at ultra-Planckian energies scales as $1/E^{2}$ in the limit $E\rightarrow\infty$, as expected for a UV-complete theory of gravity. However, the positivity of $|\mathcal{M}|^{2}$ is not manifest; it depends on the kinematical invariants, the number of colors, and the values of the gravitational couplings at the relevant energy scale. We now proceed to examine this behavior in more detail.
@@
$|\mathcal{M}|^{2}$ $=$ $-\frac{8}{3\sin^{6}(\theta)}\Big{[}f_{0}^{4}\left(25037\cos^{6}{\theta}-74631\cos^{4}{\theta}+74151\cos^{2}{\theta}-24557\right)$ (14)
$+f_{0}^{2}f_{2}^{2}\left(-269\cos^{6}{\theta}-10775\cos^{4}{\theta}+6197\cos^{2}{\theta}+4847\right)$
$+f_{2}^{4}\left(-24\cos^{8}{\theta}+216\cos^{6}{\theta}+1976\cos^{4}{\theta}+3736\cos^{2}{\theta}+240\right)\Big{]},$
@@
Assuming $f_{0}\sim f_{2}$, and disregarding the peaks near $\theta=0$ and $\theta=\pi$, which reflect the typical collinear divergences associated with gluon scattering, the squared amplitude $|\mathcal{M}|^{2}$ is positive and well-behaved in the vicinity of $\theta=\pi/2$. In fact, the positivity of $|\mathcal{M}|^{2}$ is strongly dependent on the values of the gravitational couplings.
For the case $f_{0}\sim f_{2}$, we find that $|\mathcal{M}|^{2}>0$ in the angular range $0.97\lesssim\theta\lesssim 2.18$. When $f_{2}\sim f_{0}/10$, the situation improves significantly, with $|\mathcal{M}|^{2}>0$ in the broader interval $0.31\lesssim\theta\lesssim 2.83$. In the limiting case where $f_{2}\rightarrow 0$, the amplitude becomes positive in the range $0.14\lesssim\theta\lesssim 3.00$.
@@
$|\mathcal{M}|^{2}=40f_{0}^{4}(N_{c}^{2}-1)\frac{s}{t}+\frac{f_{0}^{4}(N_{c}^{2}-1)(2805N_{c}^{2}-328)}{3}+\mathcal{O}(t).$ (15)
@@
It is straightforward to observe from Eqs. (14) and (17) that the differential cross section scales as $1/s$ in the ultra-Planckian limit. This behavior is consistent with what is expected from a well-behaved quantum theory of gravity. Notably, the result obtained for gluon scattering closely resembles that found in the case of photon-photon scattering, as discussed in Ref. Holdom:2021oii .
@@
Specializing to QCD with $N_{c}=3$ and using the Mandelstam identity $s=-t-u$, the amplitude simplifies to
@@
It is important to note that the results discussed above were obtained in the massless quark limit. In this regime, the squared amplitude exhibits an infrared divergence in the forward direction, as $t\rightarrow 0$ when $\theta\rightarrow 0$. However, the inclusion of nonzero quark masses modifies the kinematics of the scattering process, introducing a natural infrared cutoff as discussed before. In particular, the Mandelstam variable $t$ no longer vanishes in the forward limit due to the finite mass, and as a consequence, the divergence in $|\mathcal{M}|^{2}$ is regulated. This leads to a finite and well-behaved differential cross section for all scattering angles. The mass of the quark thus plays the role of an infrared regulator, softening the singular behavior of the gravitational interaction in the $t$-channel and rendering the forward scattering physically meaningful.
@@ IV Final Remarks
In this work, we have investigated several scattering processes involving quarks and gluons mediated by graviton exchange within the framework of Agravity—a scale-invariant and renormalizable theory of quadratic quantum gravity. Focusing on the ultra-Planckian regime, we computed the tree-level amplitudes for $gg\rightarrow gg$, $gg\rightarrow q\bar{q}$, $gq\rightarrow gq$, and $qq\rightarrow qq$, and analyzed the behavior of the squared amplitudes and the corresponding differential cross sections.
One of the questions addressed in this paper concerns the compatibility between unitarity and UV completeness in quadratic gravity models that inherently include ghost-like states. By isolating the gravitational contributions and examining the positivity properties of the squared amplitudes, we assessed whether unitarity-preserving behavior can persist in spite of the presence of higher-derivative ghosts. Our results indicate that the squared amplitudes remain positive in a wide region of the angular domain, particularly when the gravitational couplings satisfy $f_{2}\ll f_{0}$, as is naturally expected due to their renormalization group running in the deep UV. The analysis also shows that the differential cross sections for all studied processes scale as $1/s$ in the high-energy limit. This behavior, consistent with a Froissart-like bound, supports the interpretation of Agravity as a UV-complete framework.

### v3

@@ Abstract
We investigate tree-level scattering processes involving quarks ($q$) and gluons ($g$) mediated by graviton exchange in the framework of agravity, a dimensionless and renormalizable theory of quadratic quantum gravity. Focusing on the ultra-Planckian regime, characterized by the Mandelstam variable $s=(p_{1}+p_{2})^{2}$, which corresponds to the total energy squared in the center-of-momentum frame, being much larger than any particle mass scale, we compute the squared amplitudes and analyze the differential cross sections for the processes $g\to gg, gg$$g\to q\bar{q}$, $gq\to gq$, and $qq\to qq$. We demonstrate that all amplitudes scale as $1/s$ at high energies, in agreement with expectations for a UV-complete theory of gravity. In addition, we explore the issue of unitarity in the presence of higher-derivative ghost modes by analyzing the positivity properties of the squared amplitudes. While IR divergences appear in the forward scattering of massless particles, we show that these are regularized by finite quark masses. Our findings support the viability of agravity as a perturbatively unitary and UV-complete extension of general relativity, capable of consistently describing gravitational interactions among elementary matter fields at trans-Planckian energies.
@@ I Introduction
In this context, quadratic quantum gravity (QQG) has been extensively studied as a promising candidate for a renormalizable theory of quantum gravity Stelle:1976gc ; Tomboulis:1977jk ; Odintsov:1991nd ; Salvio:2014soa ; Salvio:2017qkx ; Einhorn:2014gfa ; Buchbinder:2017lnd ; Donoghue:2021cza . Although such theories are renormalizable, they typically suffer from issues related to the presence of ghostlike or tachyonic modes. Recent investigations Salvio:2014soa ; Alvarez-Gaume:2015rwa ; Holdom:2015kbf ; Anselmi:2018ibi ; Salvio:2018crh ; Donoghue:2021cza ; Buccio:2024hys suggest that while ghost states might be controllable under certain prescriptions, the presence of tachyonic instabilities is generally regarded as a fundamental obstacle.
@@
Although agravity is free from tachyonic instabilities, it is not, in principle, exempt from the presence of ghostlike states. It is worth emphasizing that the issue of unitarity in QQG has been examined in the recent literature. In Ref. Salvio:2019wcp , it was shown that when properly quantized, QQG can remain unitary despite the presence of higher-derivative terms. This analysis was further extended in Ref. Salvio:2024joi , where the proof of unitarity was generalized to the nonperturbative and background-independent regime, reinforcing the consistency of these theories beyond perturbation theory. In addition, IR properties of agravity were studied in Ref. Salvio:2018kwh , where the authors identified an IR enhancement in tree-level scattering amplitudes involving scalar, fermionic, and gauge fields.
While these results strongly support the theoretical consistency of agravity at both the perturbative and nonperturbative levels, it is also instructive to assess how such consistency manifests in explicit scattering processes involving matter and gauge fields. Such amplitude-based studies offer an alternative window into the interplay between higher-derivative dynamics and unitarity, providing valuable insight into the physical viability of agravity when applied to high-energy processes involving gauge and matter fields.
Progress in this direction has been made by Holdom in Refs. Holdom:2021oii ; Holdom:2021hlo . In Ref. Holdom:2021oii , the author demonstrates that QQG leads to a well behaved differential cross section for photon-photon scattering across all energy scales. The study also highlights how unitarity may be preserved even in the absence of positivity, with significant implications for the structure of scattering amplitudes and cross sections. In a subsequent work, Ref. Holdom:2021hlo , it is shown that high-energy scattering cross sections in QQG exhibit a $1/E^{2}$ scaling due to remarkable cancellations among the contributions of different graviton modes. This behavior provides further support for the idea that QQG may constitute a viable UV-complete theory of quantum gravity, capable of maintaining unitarity without relying on the positivity of individual amplitudes.
@@
The main objective of the present work is to compute graviton-mediated scattering amplitudes for quarks and gluons within the framework of QQG, specifically in the ultra-Planckian regime. To this end, we couple massless QCD fields to agravity, a dimensionless and renormalizable model of QQG, and analyze the positivity of the squared scattering amplitudes, as well as the behavior of the corresponding differential cross sections in the deep UV.
@@ III Agravity-Mediated Scattering Processes
$\mathcal{M}_{t}$ $=$ $\mathcal{M}_{s}\Big|_{s\leftrightarrow t,\penalty 10000\ p_{3}\leftrightarrow-p_{2},\penalty 10000\ b\leftrightarrow c},$ (10)
$\mathcal{M}_{u}$ $=$ $\mathcal{M}_{s}\Big|_{u\leftrightarrow t,\penalty 10000\ p_{3}\leftrightarrow p_{4},\penalty 10000\ c\leftrightarrow d}.$ (11)
@@
$|\mathcal{M}|^{2}$ $=\frac{f_{0}^{4}\,t^{2}u^{2}}{576\,s^{4}t^{2}u^{2}}\Big[3(979N_{c}^{2}-675)t^{4}+4648N_{c}^{2}t^{3}u+2(2336N_{c}^{2}+2025)t^{2}u^{2}$
$\quad+4648N_{c}^{2}tu^{3}+3(979N_{c}^{2}-675)u^{4}\Big]$
$\quad+\frac{4f_{2}^{4}}{576\,s^{4}t^{2}u^{2}}\Big[36(N_{c}^{2}-1)t^{8}+180(N_{c}^{2}-1)t^{7}u+3(205N_{c}^{2}-122)t^{6}u^{2}$
$\quad+4(244N_{c}^{2}-45)t^{5}u^{3}+2(581N_{c}^{2}+18)t^{4}u^{4}+4(244N_{c}^{2}-45)t^{3}u^{5}$
$\quad+3(205N_{c}^{2}-122)t^{2}u^{6}+180(N_{c}^{2}-1)tu^{7}+36(N_{c}^{2}-1)u^{8}\Big]$
$\quad-\frac{8\,f_{0}^{2}f_{2}^{2}\,t^{2}u^{2}}{576\,s^{4}t^{2}u^{2}}\Big[3(109N_{c}^{2}-67)t^{4}+(511N_{c}^{2}-36)t^{3}u+2(209N_{c}^{2}+315)t^{2}u^{2}$
$\quad+(511N_{c}^{2}-36)tu^{3}+3(109N_{c}^{2}-67)u^{4}\Big]\,$, (12)
where $N_{c}$ denotes the number of colors, with $N_{c}=3$ for QCD. The expression for $|\mathcal{M}|^{2}$ shown above corresponds to the squared matrix element fully summed over the color and polarization degrees of freedom of both the initial- and final-state gluons  [^1] Technical note on polarization sums. The original calculation replaced the polarization sum by $\sum_{\lambda}\epsilon_{\lambda}^{\mu}(p)\epsilon_{\lambda}^{\nu*}(p)\to-\eta^{\mu\nu}$, which neglects the dependence on the gauge-fixing condition. The correct physical polarization sum reads $\sum_{\lambda}\epsilon^{\mu}_{\lambda}(p)\epsilon^{\nu*}_{\lambda}(p)=-\eta^{\mu\nu}+\frac{p^{\mu}n^{\nu}+p^{\nu}n^{\mu}}{p\cdot n}\,$, where $n^{\mu}$ is a reference vector satisfying $n\cdot\epsilon(p)=0$. We adopt the choice $n_{1}^{\mu}=\frac{p_{2}^{\mu}}{\omega},\quad n_{2}^{\mu}=\frac{p_{1}^{\mu}}{\omega},\quad n_{3}^{\mu}=\frac{p_{4}^{\mu}}{\omega},\quad n_{4}^{\mu}=\frac{p_{3}^{\mu}}{\omega},$ valid in the CoM reference frame, where $\omega=\frac{\sqrt{s}}{2}$. This choice corresponds to the axial (physical) gauge and ensures that only the two transverse physical polarizations contribute to the scattering amplitude. It also eliminates spurious longitudinal components and leads to the correct tensor structure of the squared amplitude..
It is important to note that, similarly to the ultra-Planckian scattering of photons Holdom:2021oii , the differential cross section for gluon scattering at ultra-Planckian energies scales as $1/E^{2}$ in the limit $E\rightarrow\infty$, as expected for a UV-complete theory of gravity.
@@
$|\mathcal{M}|^{2}$ $=\frac{1}{12288\,\sin^{4}\theta}\Bigg[49152\,f_{2}^{4}\!\left(5+3\cos 2\theta\right)+\,\sin^{4}\theta\,\Big(377529f_{0}^{4}-334328f_{0}^{2}f_{2}^{2}$
$\qquad+158288f_{2}^{4}+20(7065f_{0}^{4}-6632f_{0}^{2}f_{2}^{2}+1808f_{2}^{4})\cos 2\theta$
+ $\qquad+\,5(375f_{0}^{4}-200f_{0}^{2}f_{2}^{2}+176f_{2}^{4})\cos 4\theta\Big)\Bigg]\,$, (14)
@@
The expression (III) is strictly positive for all scattering angles $0<\theta<\pi$ and for arbitrary real couplings $f_{0}$ and $f_{2}$. The collinear structure is entirely contained in the prefactor of the first term, proportional to $f_{2}^{4}$, leading to the behaviour $|{\mathcal{M}}|^{2}\sim f_{2}^{4}/\sin^{4}\theta$ near $\theta\to 0,\pi$. This behaviour is the expected enhancement from spin-2 exchange in the $t$- and $u$-channels. In the limit where $f_{2}$ becomes asymptotically small (as expected from asymptotic freedom), this enhancement is suppressed and the angular distribution remains finite for all physical scattering angles.
@@
$|\mathcal{M}|^{2}$ $=f_{2}^{4}(N_{c}^{2}-1)\left(\frac{s^{2}}{4t^{2}}+\frac{s}{4t}\right)$
$+\frac{1}{192}\Big[f_{0}^{4}\big(979\,N_{c}^{2}-675\big)+4f_{2}^{4}\big(85\,N_{c}^{2}-2\big)-8f_{0}^{2}f_{2}^{2}\big(109\,N_{c}^{2}-67\big)\Big]+\mathcal{O}(t/s).$ (15)
@@
The result (III) shows explicitly the emergence of collinear singularities as $t\to 0$, controlled by the terms proportional to $f_{2}^{4}$. If $f_{2}$ is treated as a fixed, nonvanishing parameter, the amplitude diverges as $t^{-2}$ and $t^{-1}$, as expected from massless spin-2 exchange. However, in a realistic UV scenario where $f_{2}$ is asymptotically free, these contributions become parametrically suppressed and the forward amplitude remains finite and IR safe. In this case, the scattering amplitude is well-behaved even for small scattering angles, consistent with the expected properties of a UV-complete dimensionless gravitational theory.
@@
It is straightforward to observe from Eqs. (III) and (17) that the differential cross section scales as $1/s$ in the ultra-Planckian limit. This behavior is consistent with what is expected from a well behaved quantum theory of gravity. Notably, the result obtained for gluon scattering closely resembles that found in the case of photon-photon scattering, as discussed in Ref. Holdom:2021oii .
@@
Specializing to QCD with $N_{c}=3$ and using the Mandelstam identity s=tus=-t-u, the amplitude simplifies to
@@
It is important to note that the results discussed above were obtained in the massless quark limit. In this regime, the squared amplitude exhibits an infrared divergence in the forward direction, as $t\rightarrow 0$ when $\theta\rightarrow 0$. However, the inclusion of nonzero quark masses modifies the kinematics of the scattering process, introducing a natural infrared cutoff as discussed before. In particular, the Mandelstam variable $t$ no longer vanishes in the forward limit due to the finite mass, and as a consequence, the divergence in $|\mathcal{M}|^{2}$ is regulated. This leads to a finite and well behaved differential cross section for all scattering angles. The mass of the quark thus plays the role of an infrared regulator, softening the singular behavior of the gravitational interaction in the $t$-channel and rendering the forward scattering physically meaningful.
@@ IV Final Remarks
In this work, we have investigated several scattering processes involving quarks and gluons mediated by graviton exchange within the framework of agravity—a scale-invariant and renormalizable theory of quadratic quantum gravity. Focusing on the ultra-Planckian regime, we computed the tree-level amplitudes for $g\rightarrow gg, gg$$g\rightarrow q\bar{q}$, $gq\rightarrow gq$, and $qq\rightarrow qq$, and analyzed the behavior of the squared amplitudes and the corresponding differential cross sections.
One of the questions addressed in this paper concerns the compatibility between unitarity and UV completeness in quadratic gravity models that inherently include ghostlike states. By isolating the gravitational contributions and examining the positivity properties of the squared amplitudes, we assessed whether unitarity-preserving behavior can persist in spite of the presence of higher-derivative ghosts. Our results indicate that the squared amplitudes remain positive in a wide region of the angular domain, particularly when the gravitational couplings satisfy $f_{2}\ll f_{0}$, as is naturally expected due to their renormalization group running in the deep UV. The analysis also shows that the differential cross sections for all studied processes scale as $1/s$ in the high-energy limit. This behavior, consistent with a Froissart-like bound, supports the interpretation of agravity as a UV-complete framework.